a uniform metre rule of mass 150g is balanced on a knife edge of xm from a mass of 50g suspended at the 0.1m mark. find the position (x) of the knife edge
1
Expert's answer
2021-09-19T09:25:35-0400

Given:

x - the mark on the ruler where the knife edge is

M - mass of the ruler

m - mass of the object attached to the ruler

a - the point where the mass m is attached

L - length of the ruler

Solution

The torque created by the hanging mass and the piece of ruler that is on the side where the mass is attached:


"\tau_1=\frac{x}{L}M\frac{x}{2}+am."

The torque created by the ruler mass on the other side:


"\tau_2=\frac{L-x}{L}m\frac{L-x}{2}."

Equilibrium means that


"\tau_1=\tau_2,\\\space\\
\frac{x^2}{2L}M+am=\frac{(L-x)^2}{2L}M,\\\space\\
x=0.467\text{ m}."